OpenStudy (wampominater):

Verify the identity. quantity one minus sine of x divided by cosine of x equals cosine of x divided by quantity one plus sine of x

OpenStudy (wampominater):

looks like this \[\frac{ 1 - sinx }{ cosx } = \frac{ cosx }{ 1 + sinx }\]

Nnesha (nnesha):

cross multiply ?

Nnesha (nnesha):

or i would solve right to make it equal

Nnesha (nnesha):

right side*

OpenStudy (wampominater):

ok. so how would I start with right side?

Nnesha (nnesha):

you need to multiply denominator and numerator by the conjugate of 1+sinx

OpenStudy (wampominater):

ahh ok

OpenStudy (wampominater):

so it is \[\frac{ cosx - sinxcosx }{ 1-\sin^2x}\]

Nnesha (nnesha):

okay so you don't need to distribute 1-sinx by cosx at the numerator \[\frac{ \cos(x)(1-\sin(x)) }{ 1-\sin^2 }\] apply the identity 1-sin^2x equal to what ?

OpenStudy (wampominater):

cos^2x

OpenStudy (wampominater):

so it would be \[\frac{ cosx(1-sinx) }{ \cos^2x }\]

Nnesha (nnesha):

yes right cos^2x can be written as cos x times cos x \[\huge\rm \frac{ \cos(x)(1-\sin(x)) }{ cos(x) \times cos(x) }\]

OpenStudy (wampominater):

so factor out top and bottom leaves me with \[\frac{ 1-sinx }{ cosx }\]

Nnesha (nnesha):

yep

OpenStudy (wampominater):

and thats it

OpenStudy (wampominater):

ty!

Nnesha (nnesha):

yep

Nnesha (nnesha):

my pleasure!